OpenStudy (anonymous):

A clown weighs 60 lb more than a trapeze artist. The trapeze artist weighs two thirds as much of the clown. How much does each weigh? Choices for the weight of the trapeze artist: 110, 120,125.

OpenStudy (anonymous):

Ok lets start with an expression. : c = t + 60 --------- The trapeze artist weighs 1/3's as much as the clown. \[t =\frac{ 1 }{ 3 }c\]

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

Ok. Replace c in the fraction for t+60. \[t =\frac{ 1 }{ 3 }t + 60\] -\[t = \frac{ 2 }{ 3 } + \frac{ 2 }{ 3 }\times 60\]

OpenStudy (anonymous):

It looks complicated but it's not. The answer would be 180.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You're welcome.